# 3HDM at the ILC

### v1

@@ Abstract
The Three-Higgs-Doublet Model (3HDM) extends the Standard Model by introducing two additional scalar doublets, leading to a rich spectrum of new particles: three neutral CP-even Higgs bosons ($h_{1}$, $H_{2}$, $H_{3}$), two neutral CP-odd Higgs bosons ($A_{2}$, $A_{3}$), and two charged Higgs bosons ($H_{2}^{+}$, $H_{3}^{+}$). In this work, we present a phenomenological study of the 3HDM at the future International Linear Collider (ILC) with a center-of-mass energy of $\sqrt{s}=1000\,\text{GeV}$. Applying a comprehensive set of theoretical and experimental constraints, we identify promising new physics signals with sufficiently large production cross-sections. Our analysis shows that $e^{+}e^{-}\to H_{2}A_{2}$, $e^{+}e^{-}\to H_{2}H_{2}Z$, $A_{2}A_{2}Z$, $H_{2}H_{2}^{\pm}W^{\mp}$, $A_{2}H_{2}^{\pm}W^{\mp}$ and $H_{1}H_{2}A_{2}$ are among the most sensitive channels to probe this extended Higgs sector. We demonstrate that a future ILC would offer a powerful platform to test these interactions and discover these heavier Higgs bosons thus providing evidence of physics beyond the Standard Model.
@@ I Introduction
The Standard Model (SM) of particle physics has been the most successful theory so far, verified and established by past and current experimental setups. The final validation of the SM came with the discovery of the Higgs boson at the Large Hadron Collider (LHC) in 2012 [1, 2]. Although the SM has been the cornerstone of particle physics, it has long been known that it needs to be supplanted with new dynamics at the TeV scale owing to many theoretical and experimental reasons - the evidence of dark matter and neutrino mass are two of the most pressing theoretical issues that do not find a resolution within the SM. One of the many ways the theoretical framework of the SM can be extended is by adding additional scalar sectors, such as singlet [3, 4] and doublet extensions [5], and their generalization to multi-Higgs doublet models [6, 7]. Of late, along these lines, the 3HDM has emerged as a potentially interesting and promising source of investigation [8, 9]. These models have been key in addressing unresolved questions, such as providing new sources of CP violation which are crucial for explaining baryogenesis [10], enabling strong first-order electroweak phase transitions, a necessary condition for successful electroweak baryogenesis [11]. In certain realizations, the extended scalar sector can also offer viable dark matter candidates through neutral scalars [12]. Beyond cosmological implications, such models predict a richer Higgs spectrum with distinct collider signatures, offering promising discovery opportunities at current and future colliders like the LHC and ILC. While models with two Higgs doublets have been extensively studied [5], the 3HDM offers an even richer framework for addressing fundamental questions beyond the Standard Model. Furthermore, the additional scalar degrees of freedom enable more flexible Yukawa structures, providing potential explanations for the observed patterns of fermion masses and mixing [13]. From a phenomenological perspective, the extended scalar spectrum enables a variety of novel collider signatures including multiple charged and neutral Higgs bosons that can be probed at future experiments such as the International Linear Collider (ILC) or the Future Circular Collider (FCC-ee).
@@
In this paper, we explore the phenomenological aspects of the 3HDM from the standpoint of the future ILC experiment. As the ILC global project proposal [14, 15] continues to evolve and gain traction within the community, it is timely and important to investigate the avenues through which various Beyond Standard Model (BSM) scenarios could manifest in its setup. With this motivation, we investigate the phenomenology of the heavy Higgs bosons in the 3HDM at the ILC operating at a center-of-mass energy of $\sqrt{s}=$ 1000 GeV [16]. We focus on channels that are expected to have reasonable cross sections while operating with benchmark points allowed by all existing theoretical and experimental constraints. The paper is organized as follows: in Sec. II, we provide a brief overview of the 3HDM concentraing on the scalar structure, and list the theoretical and experimental constraints on the model. The phenomenological analysis of the various relevant channels at the ILC is presented in Sec. III, and we present our conclusions in Sec. IV.
@@ II.1 Scalar Sector
where $k=1,2,3$. We consider a $Z_{3}$ symmetric potential under which the Higgs fields transform as
@@
where $\omega=e^{2\pi i/3}$ are the cube roots of unity. The most general $SU(2)_{L}\times U(1)_{Y}$ invariant potential preserving the $Z_{3}$ symmetry is given by
$\begin{split}V&=m_{11}^{2}(\Phi_{1}^{\dagger}\Phi_{1})+m_{22}^{2}(\Phi_{2}^{\dagger}\Phi_{2})+m_{33}^{2}(\Phi_{3}^{\dagger}\Phi_{3})\\ &+\lambda_{1}(\Phi_{1}^{\dagger}\Phi_{1})^{2}+\lambda_{2}(\Phi_{2}^{\dagger}\Phi_{2})^{2}+\lambda_{3}(\Phi_{3}^{\dagger}\Phi_{3})^{2}\\ &+\lambda_{4}(\Phi_{1}^{\dagger}\Phi_{1})(\Phi_{2}^{\dagger}\Phi_{2})+\lambda_{5}(\Phi_{1}^{\dagger}\Phi_{1})(\Phi_{3}^{\dagger}\Phi_{3})+\lambda_{6}(\Phi_{2}^{\dagger}\Phi_{2})(\Phi_{3}^{\dagger}\Phi_{3})\\ &+\lambda_{7}(\Phi_{1}^{\dagger}\Phi_{2})(\Phi_{2}^{\dagger}\Phi_{1})+\lambda_{8}(\Phi_{1}^{\dagger}\Phi_{3})(\Phi_{3}^{\dagger}\Phi_{1})+\lambda_{9}(\Phi_{2}^{\dagger}\Phi_{3})(\Phi_{3}^{\dagger}\Phi_{2})\\ &+[\lambda_{10}(\Phi_{1}^{\dagger}\Phi_{2})(\Phi_{1}^{\dagger}\Phi_{3})+\lambda_{11}(\Phi_{1}^{\dagger}\Phi_{2})(\Phi_{3}^{\dagger}\Phi_{2})+\lambda_{12}(\Phi_{1}^{\dagger}\Phi_{3})(\Phi_{2}^{\dagger}\Phi_{3})+h.c.].\\ &\end{split}$ (3)
Here, in general $\lambda_{1,...,9}$ are real parameters, while $\lambda_{10},\lambda_{11},\lambda_{12}$ can be complex. However, here we disregard the complex phases of the potential to prevent mixing between the CP-odd and CP-even states [8]. The relevant mass terms for the CP-even Higgs are deduced from the scalar potential in Eqn. 3, and can be written as
@@
$O_{\alpha}.\mathcal{M}^{2}_{S}.O_{\alpha}^{T}=\begin{pmatrix}m_{H1}^{2}&0&0\\ 0&m_{H2}^{2}&0\\ 0&0&m_{H3}^{2}\end{pmatrix}$ (4)
@@
The mass terms for the charged Higgses can similarly be extracted from the scalar potential given in Eqn. 3:
@@
$(O_{\gamma 2}.O_{\beta})^{\dagger}=\begin{pmatrix}c_{\beta 1}c_{\beta 2}&-c_{\gamma 2}s_{\beta 1}+c_{\beta 1}s_{\beta 2}s_{\gamma 2}&-c_{\beta 1}c_{\gamma 2}s_{\beta 2}-s_{\beta 1}s_{\gamma 2}\\ c_{\beta 2}s_{\beta 1}&c_{\beta 1}c_{\gamma 2}+s_{\beta 1}s_{\beta 2}s_{\gamma 2}&-c_{\gamma 2}s_{\beta 1}s_{\beta 2}+c_{\beta 1}s_{\gamma 2}\\ s_{\beta 2}&-c_{\beta 2}s_{\gamma 2}&c_{\beta 2}c_{\gamma 2}\end{pmatrix}.$ (6)
@@ II.2 Yukawa Sector
Experimental limits place stringent constraints on the presence of tree-level Flavor Changing Neutral Currents (FCNCs), posing a significant challenge for multi-Higgs models. To address this, we impose Natural Flavor Conservation (NFC), a condition that restricts each Higgs doublet to couple with only one class of fermions. In the framework we adopt for the present work, we employ the Type-Z or democratic Yukawa structure, wherein the masses of the leptons, the down-type quarks, and the up-type quarks are generated by the $\Phi_{1}$, $\Phi_{2}$, and $\Phi_{3}$ doublets respectively. This assignment not only satisfies the NFC criterion but also provides a natural mechanism to account for the observed hierarchy in fermion masses. We can thus write the Yukawa Lagrangian as
@@
We work with a $Z_{3}$ symmetric potential as given in Eqn. 2 - for the Yukawa Lagrangian to remain invariant under the same, the right handed fermion fields transform as
@@ II.3 Alignment-Limit
  • $k=1\implies\alpha_{1}=\beta_{1}+2n\pi$ and $\alpha_{2}=\beta_{2}+2n\pi$.
The case with $k=-1$ is equivalent to the current scenario up to an inversion, whereas the case with $k\neq 1$ represents a special configuration in which the entire vacuum expectation value (vev) is confined to the third doublet. Therefore, we proceed with our analysis by choosing $k=1$ [8].
@@ II.4 Constraints
    Vacuum Stability: To ensure stability, the scalar potential must be bounded from below in all directions of the field space. This translates into a set of bounds on the scalar coupling parameters $\lambda_{i}$ which are listed below,
@@
    Unitarity: The unitarity constraints are necessary to ensure that the theory remains predictive at high energies. Using the results of [17, 18] for unitarity Constraints in the $Z_{3}$ symmetric 3HDM, the condition to be imposed on the $21$ eigenvalues $\Lambda_{i}$ of the relevant scattering matrices is
@@ III.1 Relevant Production Modes
We begin by listing all relevant production channels that can potentially uncover signals of a heavy Higgs in $e^{+}e^{-}$ collisions in Table 1. For each production mode, we have considered two sets of benchmark mass points, denoted as $BP\text{-}X$ and $BP\text{-}Y$, where $X$ and $Y$ are numerical labels assigned to clearly distinguish between different benchmark points across all production modes. These benchmarks satisfy all applied constraints and are thus suitable for a detailed analysis. Depending on the production mode, we consider either fully hadronic decays (FHD) or semi-leptonic decays (SLD) - the corresponding cross sections, branching ratios for the relevant signal channels, and the associated background processes will be discussed in the upcoming sections dedicated to the detailed phenomenological analysis.
@@
$SS$ $H_{2}A_{2}$ (III.2) $4b$ $BP\text{-}1$ , $BP\text{-}2$ : Table 2 $(FHD)$
$SSS$ $h_{1}H_{2}A_{2}$ (III.3) $6b$ $BP\text{-}3$ , $BP\text{-}4$ : Table 5 $(FHD)$
$SSV$ $H_{2}H_{2}Z$ (III.4) $4b+2j$ $BP\text{-}5$ , $BP\text{-}6$ : Table 9 $(FHD)$
$A_{2}A_{2}Z$ (III.5) $4b+2j$ $BP\text{-}7$ , $BP\text{-}8$ : Table 11 $(FHD)$
$H_{2}H_{2}Z$ (III.6) $4b+2\ell$ $BP\text{-}9$ , $BP\text{-}10$ : Table 14 $(SLD)$
$A_{2}A_{2}Z$ (III.7) $4b+2\ell$ $BP\text{-}11$ , $BP\text{-}12$ : Table 16 $(SLD)$
$H_{2}H_{2}^{\pm}W^{\mp}$ (III.8) $4b+4j$ $BP\text{-}13$ , $BP\text{-}14$ : Table 19 $(FHD)$
$A_{2}H_{2}^{\pm}W^{\mp}$ (III.9) $4b+4j$ $BP\text{-}15$ , $BP\text{-}16$ : Table 21 $(FHD)$
@@
The 3HDM, as mentioned in Sec. II, has multiple neutral and charged Higgs bosons. The choice of the specific channels in the table above is dictated by the parameter space that is allowed after all constraints have been imposed, and also demanding that the relevant couplings - and the associated cross-sections and branching ratios - be large enough that one can in principle probe them at colliders. This also explains why we have chosen a varied set of benchmark points.
@@ III.2 $H_{2}A_{2}$ Production: $FHD$ Mode
For the $H_{2}A_{2}$ production mode, we consider the fully hadronic decay channel which in this case means that both $H_{2}$ and $A_{2}$ decay to $b\bar{b}$ resulting in a $4b$ final state. As previously mentioned, we study the efficacy of this channel (like all channels to follow) with two sets benchmark points ($BP$), with the corresponding branching ratios ($BR$) and total cross sections $(\sigma)$ provided in Table 2. In Fig. 1, we display a few illustrative Feynman diagrams for the production process. For the purpose of phenomenological analysis, we also identify the most relevant background processes—both reducible and irreducible—which are listed in Table 3.
@@
$BP-1$:Particles $\rightarrow$ $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}A_{2}$ $H_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ Total $\sigma$ (pb)
Mass (GeV) 125 340.46 409.61 274.85 469.73 376.15 400.29 $\sigma||BR$ 0.00598 0.75197 0.70063 0.0031528
$BP-2$:Particles $\rightarrow$ $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}A_{2}$ $H_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ Total $\sigma$ (pb)
Mass (GeV) 125 287.95 457.92 209.50 425.93 330.31 441.07 $\sigma||BR$ 0.00782 0.74378 0.64988 0.0037801
Table 2: The upper portion of the table corresponds to the $BP\text{-}1$ mass points, $BR$ and total cross section ($\sigma$), while the lower portion corresponds to $BP\text{-}2$, for the $H_{2}A_{2}$ production mode and its fully hadronic decay.
Background Process Final State Cross Section($\sigma$) (pb)
@@
Table 3: Relevant Standard Model background processes listed for $e^{+}e^{-}\to H_{2}A_{2}$, with $H_{2}\to b\bar{b}$ and $A_{2}\to b\bar{b}$.
[…]
Figure 2: The plots in the upper panel represent the selection cuts applied to $BP$-1, while those in the lower panel correspond to $BP$-2. Both sets of plots show the selection cuts on $p_{T}(b_{1})$ and $p_{T}(b_{2})$. The shaded regions indicate the selected ranges that are implemented in the cutflow chart.
@@
$BP-1$ : Cuts Signal (S) Background (B) S vs B
@@
$BP-2$ : Cuts Signal (S) Background (B) S vs B
Initial (no cut) 378 7509
$N(b)=4$ 64.79 85.11 5.292
SEL: $P_{T}(b_{1})>200$ 59.72 48.24 5.747
SEL: $P_{T}(b_{2})>150$ 54.81 35.2 5.779
Table 4: The upper table represents the cutflow chart for $BP-1$, while the lower table corresponds to $BP-2$. Both cases are considered at an integrated luminosity of $100\,\mathrm{fb}^{-1}$ in the analysis. In this setup, a discovery potential of $5\sigma$ is anticipated for the signal.

### v2

# ILC Phenomenology of the $Z_{3}$ symmetric Type-Z Three Higgs Doublet Model

@@ Abstract
The Three-Higgs-Doublet Model (3HDM) extends the Standard Model by introducing two additional scalar doublets, leading to a rich spectrum of new particles: three neutral CP-even Higgs bosons ($h_{1}$, $H_{2}$, $H_{3}$), two neutral CP-odd Higgs bosons ($A_{2}$, $A_{3}$), and two charged Higgs bosons ($H_{2}^{+}$, $H_{3}^{+}$). In this work, we present a phenomenological study of the 3HDM at the future International Linear Collider (ILC) with a center-of-mass energy of $\sqrt{s}=1000\,\text{GeV}$. Applying a comprehensive set of theoretical and experimental constraints, we identify promising new physics signals with sufficiently large production cross-sections. Our analysis shows that $^{+}e^{-}\to H_{2}A_{2}, e+e$$^{+}e^{-}\to H_{2}H_{2}Z$, $A_{2}A_{2}Z$, $H_{2}H_{2}^{\pm}W^{\mp}$, $A_{2}H_{2}^{\pm}W^{\mp}$ and $H_{1}H_{2}A_{2}$ are among the most sensitive channels to probe this extended Higgs sector. We demonstrate that a future ILC would offer a powerful platform to test these interactions and discover these heavier Higgs bosons thus providing evidence of physics beyond the Standard Model.
@@ I Introduction
The Standard Model (SM) of particle physics has been the most successful theory so far, verified and established by past and current experimental setups. The final validation of the SM came with the discovery of the Higgs boson at the Large Hadron Collider (LHC) in 2012 [1, 2]. Although the SM has been the cornerstone of particle physics, it has long been known that it needs to be supplanted with new dynamics at the TeV scale owing to many theoretical and experimental reasons - the evidence of dark matter and neutrino mass are two of the most pressing theoretical issues that do not find a resolution within the SM. One of the many ways the theoretical framework of the SM can be extended is by adding additional scalar sectors, such as singlet [3, 4] and doublet extensions [5], and their generalization to multi-Higgs doublet models [6, 7]. Along these lines, the 3HDM [8, 9, 10, 11, 12, 13, 14, 15, 16, 17] has also been a long-standing source of investigation. These models have been key in addressing unresolved questions, such as providing new sources of CP violation which are crucial for explaining baryogenesis [18], enabling strong first-order electroweak phase transitions, a necessary condition for successful electroweak baryogenesis [19]. In certain realizations, the extended scalar sector can also offer viable dark matter candidates through neutral scalars [20]. Beyond cosmological implications, such models predict a richer Higgs spectrum with distinct collider signatures, offering promising discovery opportunities at current and future colliders like the LHC and International Linear Collider (ILC) [21]. While models with two Higgs doublets have been extensively studied [5], the 3HDM offers an even richer framework for addressing fundamental questions beyond the Standard Model. Furthermore, the additional scalar degrees of freedom enable more flexible Yukawa structures, providing potential explanations for the observed patterns of fermion masses and mixing [22]. From a phenomenological perspective, the extended scalar spectrum enables a variety of novel collider signatures including multiple charged and neutral Higgs bosons that can be probed at future experiments such as the ILC or the Future Circular Collider (FCC-ee) [23]. For a complete review of BSM scenarios with extended scalar sectors, the reader is invited to consult Ref. [13].
@@
In this paper, we explore the phenomenological aspects of the $Z_{3}$ symmetric Type Z 3HDM from the standpoint of the future ILC experiment. As the ILC global project proposal [24, 25] continues to evolve and gain traction within the community, it is timely and important to investigate the avenues through which various Beyond Standard Model (BSM) scenarios could manifest in its setup. With this motivation, we investigate the phenomenology of the heavy Higgs bosons in the 3HDM at the ILC operating at a center-of-mass energy of $\sqrt{s}=$ 1000 GeV [26]. We focus on channels that are expected to have reasonable cross sections while operating with benchmark points allowed by all existing theoretical and experimental constraints. The paper is organized as follows: in section II, we provide a brief overview of the 3HDM concentraing on the scalar structure, and list the theoretical and experimental constraints on the model. The phenomenological analysis of the various relevant channels at the ILC is presented in section III, and we present our conclusions in section IV.
@@ II.1 Scalar Sector
where $k=1,2,3$. We consider a $Z_{3}$ symmetric potential [27, 28, 29] under which the Higgs fields transform as
@@
where $\omega=e^{2\pi i/3}$ are the nontrivial cubic roots of unity. The most general $SU(2)_{L}\times U(1)_{Y}$ invariant potential preserving the $Z_{3}$ symmetry is given by
[…]
Here, in general $\lambda_{1,...,9}$ are real parameters, while $\lambda_{10},\lambda_{11},\lambda_{12}$ can be complex. However, here we disregard the complex phases of the potential to prevent mixing between the CP-odd and CP-even states [8]. Consequently, the scalar potential is CP conserving, with no explicit CP violation. Further, in what follows, we also assume there is no spontaneous CP violation as well, i.e., we assume no relative phases in the vacuum expectation values are generated through minimization. The relevant mass terms for the CP-even Higgs are deduced from the scalar potential in Equation 3, and can be written as
@@
$O_{\alpha}\mathcal{M}^{2}_{S}O_{\alpha}^{T}=\begin{pmatrix}m_{H1}^{2}&0&0\\ 0&m_{H2}^{2}&0\\ 0&0&m_{H3}^{2}\end{pmatrix}$ (4)
@@
The mass terms for the charged Higgses can similarly be extracted from the scalar potential given in Equation 3:
@@
$(O_{\gamma 2}O_{\beta})^{\dagger}=\begin{pmatrix}c_{\beta 1}c_{\beta 2}&-c_{\gamma 2}s_{\beta 1}+c_{\beta 1}s_{\beta 2}s_{\gamma 2}&-c_{\beta 1}c_{\gamma 2}s_{\beta 2}-s_{\beta 1}s_{\gamma 2}\\ c_{\beta 2}s_{\beta 1}&c_{\beta 1}c_{\gamma 2}+s_{\beta 1}s_{\beta 2}s_{\gamma 2}&-c_{\gamma 2}s_{\beta 1}s_{\beta 2}+c_{\beta 1}s_{\gamma 2}\\ s_{\beta 2}&-c_{\beta 2}s_{\gamma 2}&c_{\beta 2}c_{\gamma 2}\end{pmatrix}$. (6)
@@ II.2 Yukawa Sector
Experimental limits place stringent constraints on the presence of tree-level Flavor Changing Neutral Currents (FCNCs), posing a significant challenge for multi-Higgs models. To address this, we impose Natural Flavor Conservation (NFC), a condition that restricts each Higgs doublet to couple with only one class of fermions. In the framework we adopt for the present work, we employ the Type-Z or democratic Yukawa structure, wherein the masses of the leptons, the down-type quarks, and the up-type quarks are generated by the $\Phi_{1}$, $\Phi_{2}$, and $\Phi_{3}$ doublets respectively. We can thus write the Yukawa Lagrangian as
@@
We work with a $Z_{3}$ symmetric potential as given in Equation 2 - for the Yukawa Lagrangian to remain invariant under the same, the right handed fermion fields transform as
@@ II.3 Alignment-Limit
  • $k=1\implies\alpha_{1}=\beta_{1}$ and $\alpha_{2}=\beta_{2}$.
The case with k=1k=-1 is equivalent to the current scenario up to an inversion, whereas the case with $k\neq 1$ represents a special configuration in which the entire vacuum expectation value (vev) is confined to the third doublet. Therefore, we proceed with our analysis by choosing $k=1$ [8].
@@ II.4 Constraints
    Vacuum Stability: To ensure stability, the scalar potential must be bounded from below in all directions of the field space. [^1] We remark here that the complete necessary and sufficient conditions for the $Z_{3}$ symmetric 3HDM are not yet known. The conditions listed in Eq. 13 should instead be understood as a set of sufficient stability conditions obtained by requiring positivity of the quartic potential along a representative set of field directions in the scalar field space. These conditions are analogous to the well-known copositivity-inspired constraints [30], used in multi-Higgs models when the full analytic BFB solution is unavailable. While they do not exhaust all possible directions in field space, they guarantee stability along all tested directions and therefore ensure that the scalar potential is bounded from below in a physically relevant region of parameter space. Also see Ref [28, 31]. This translates into a set of bounds on the scalar coupling parameters $\lambda_{i}$ which are listed below,
@@
    Unitarity: The unitarity constraints are necessary to ensure that the theory remains predictive at high energies. Using the results of [32, 33] for unitarity Constraints in the $Z_{3}$ symmetric 3HDM, the condition to be imposed on the 2121 eigenvalues $\Lambda_{i}$ of the relevant scattering matrices is
@@ III.1 Relevant Production Modes
We begin by listing all relevant production channels that can potentially uncover signals of a heavy Higgs in $e^{+}e^{-}$ collisions in Table 1. For each production mode, we have considered two sets of benchmark mass points, denoted as BPX and BPY, where X and Y are numerical labels assigned to clearly distinguish between different benchmark points across all production modes. These benchmarks satisfy all applied constraints and are thus suitable for a detailed analysis. Depending on the production mode, we consider either fully hadronic decays (FHD) or semi-leptonic decays (SLD) - the corresponding cross sections, branching ratios for the relevant signal channels, and the associated background processes will be discussed in the upcoming sections dedicated to the detailed phenomenological analysis.
@@
SSSS $H_{2}A_{2}$ (III.2) $4b$ BP1 : Table 2 $(FHD)$
SSSSSS $h_{1}H_{2}A_{2}$ (III.3) $6b$ BP2 : Table 5 $(FHD)$
SSVSSV $H_{2}H_{2}Z$ (III.4) $4b+2j$ BP3 : Table 9 $(FHD)$
$A_{2}A_{2}Z$ (III.5) $4b+2j$ BP4 : Table 11 $(FHD)$
$H_{2}H_{2}Z$ (III.6) $4b+2\ell$ BP5 : Table 14 $(SLD)$
$A_{2}A_{2}Z$ (III.7) $4b+2\ell$ BP6 : Table 16 $(SLD)$
$H_{2}H_{2}^{\pm}W^{\mp}$ (III.8) $4b+4j$ BP7 : Table 19 $(FHD)$
$A_{2}H_{2}^{\pm}W^{\mp}$ (III.9) $4b+4j$ BP8 : Table 20 $(FHD)$
@@
The 3HDM, as mentioned in section II, has multiple neutral and charged Higgs bosons. The choice of the specific channels in the table above is dictated by the parameter space that is allowed after all constraints have been imposed, and also demanding that the relevant couplings - and the associated cross-sections and branching ratios - be large enough that one can in principle probe them at colliders. This also explains why we have chosen a varied set of benchmark points.
@@ III.2 $H_{2}A_{2}$ Production: $FHD$ Mode
For the $H_{2}A_{2}$ production mode, we consider the fully hadronic decay channel which in this case means that both $H_{2}$ and $A_{2}$ decay to $b\bar{b}$ resulting in a $4b$ final state. As previously mentioned, we study the efficacy of this channel (like all channels to follow) with a benchmark point (BP), with the corresponding branching ratios (BR) and total cross sections $(\sigma)$ provided in Table 2. In Figure 1, we display a few illustrative Feynman diagrams for the production process. For the purpose of phenomenological analysis, we also identify the most relevant background processes—both reducible and irreducible—which are listed in Table 3.
@@
BP1:Particles $\rightarrow$ $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}A_{2}$ $H_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ Total $\sigma$ (pb)
Mass (GeV) 125 340.46 409.61 274.85 469.73 376.15 400.29 $\sigma\times BR$ 0.00598 0.75197 0.70063 0.0031528
Table 2: The table corresponds to the BP1 mass points, BR and total cross section ($\sigma$), for the $H_{2}A_{2}$ production mode and its fully hadronic decay.
Background Process Final State Cross Section ($\sigma$) [pb]
@@
Table 3: Relevant Standard Model background processes for $e^{+}e^{-}\to H_{2}A_{2}$, with $H_{2}\to b\bar{b}$ and $A_{2}\to b\bar{b}$.
To isolate the signal in our analysis, we impose a stringent pre-selection cut of N(b)=4N(b)=4222An alternative strategy could involve a more inclusive requirement such as $N(b)\geq 4$. In realistic detector environments, additional jets may arise due to initial-state radiation (ISR), final-state radiation (FSR), and jet fragmentation. Such effects can result in a higher jet multiplicity than the parton-level expectation. Therefore, using $N(b)\geq 4$ may be beneficial in scenarios where a strict cut significantly reduces signal efficiency. However, in the present case, this issue is not severe, especially when weighed against the benefit of reduced background efficiency, and the exact $N(b)=4$ selection is sufficient and effective., thereby selecting only events that reconstruct exactly four $b$-tagged jets. Following this pre-selection, we apply additional cuts on the transverse momenta of the leading and subleading $b$-jets, $P_{T}(b_{1})$ and $P_{T}(b_{2})$, for both BP1 and BP2 as both these kinematic variables show sufficient promise in isolating signal from background as can be seen from Figure 2.
Figure 2: The plots in the upper panel represent the selection cuts applied to BP1. Both sets of plots show the selection cuts on $p_{T}(b_{1})$ and $p_{T}(b_{2})$. The shaded regions indicate the selected ranges that are implemented in the cutflow chart.
@@
BP1 : Cuts Signal (S) Background (B) S vs B
@@
Table 4: The table represents the cutflow chart for BP1. This is considered at an integrated luminosity of $100\,\mathrm{fb}^{-1}$ in the analysis. In this setup, a discovery potential of $5\sigma$ is anticipated for the signal.